# Effects of quadratic optomechanical coupling on bipartite entanglements, mechanical ground-state cooling and squeezing in an electro-optomechanical system

### v1

@@ II.3 Stationary quantum fluctuations
In Fig. 9, we have plotted the effective mean phonon number of the MO versus the ratio $g_{2}/g_{1}$ when $\Delta_{0w}=\Omega_{m}$. It is seen that in the presence of the QOC with positive sign, the effective mean number of phonons goes to a minimum of about $1.68\times 10^{-3}$ at $g_{2}/g_{1}=+8\times 10^{-3}$ [see the right inset of Fig. 9] and the MO is cooled down to its ground state. On the other hand, for negative values of the QOC the mean number of phonons increases sharply at the outset with an increase in the absolute value $\mid g_{2}/g_{1}\mid$ , reaching the maximum value $0.24$ at $g_{2}/g_{1}=-2\times 10^{-3}$, and after that it drops to a minimum of about $1.70\times 10^{-3}$ at $g_{2}/g_{1}=-0.013$ [see the left inset of Fig. 9]. Hence, the positive values of QOC can provide a better cooling performance as compared to its negative values. In Fig. 10, we have numerically plotted the effective mean phonon number of the MO as a function of the dimensionless detuning of the microwave cavity $\Delta_{0w}/\Omega_{m}$, for various bath temperatures of the MO. The left panels in the figure correspond to the case of absence of QOC ($g_{2}/g_{1}=0$), while the right panels correspond to its presence with $g_{2}/g_{1}=+8\times 10^{-3}$ for which the optimal ground state cooling of the MO is achieved [see Fig. 9]. As is seen, in the absence of QOC [Figs. 10(a) - 10(d)], when $\Delta_{0w}=\Omega_{m}$ the effective mean number of phonons of the mechanical mode goes to a minimum whose value increases with increasing the mechanical bath temperature $T$, as expected. On the other hand, we find that in the presence of QOC [Figs. 10(e) - 10(h)], the minimum attainable value of the effective number of phonons is reduced by 1 order of magnitude as compared to that for the case of absence of QOC. In fact, at each temperature of the mechanical bath there exists a certain range of $\Delta_{0w}/\Omega_{m}$, which we refer to it as a mechanical cooling window, over which the QOC gives rise to the reduction of $n_{\rm eff}$. As understood from Figs. 10(e),  10(f), and  10(g), at environment temperatures lower than $T=0.1$mK, the presence of QOC ensures that the MO is cooled down to its ground state. Interestingly, Fig. 10(h)) reveals that even at the sub-Kelvin temperature of $T=0.3$mK the mean number of phonons is less than unity, which implies that the MO is near its ground state, while this is not the case in the absence of QOC [see Fig. 10(d)].
Figure 12: (Color online)(a) Density plot of $S_{Q}$ vs $g_{2}/g_{1}$ and the normalized bare dimensionless detuning of the microwave cavity $\Delta_{0w}/\Omega_{m}$, and (b) plot of $S_{Q}$ vs $g_{2}/g_{1}$ for $\Delta_{0w}=0$. The shadowed pink region corresponds to squeezing beyond the 3 dB limit. The other parameters are the same as in Fig. 2.
It should be noted that the condition $n_{\rm eff}<1$ is not the only requirement for the ground-state cooling of the mechanical mode. In order to achieve the quantum ground-state cooling it is also necessary that both momentum and displacement variances tend to $\langle\delta\hat{P}^{2}\rangle\approx\langle\delta\hat{Q}^{2}\rangle\approx 1/2$ that is the energy equipartition should be satisfied in the optimal regime close to the ground state. Otherwise, the steady state of the system is not a thermal equilibrium state, and thus one cannot define an unambiguously effective temperature for the MO in this case genes20081 . We have examined this second requirement in Fig. 11 where we have plotted the effective mean phonon number $n_{\rm eff}$ (on logarithmic scale) and the variances $\langle\delta\hat{Q}^{2}\rangle-3/4$ and $\langle\delta\hat{P}^{2}\rangle-3/4$ versus $\Delta_{0w}/\Omega_{m}$ for $g_{2}=0$ (panels a-d) and $g_{2}=8\times 10^{-3}g_{1}$ (panels e-h), and for different values of temperature $T=1\rm mk$, $10\rm mK$, $100\rm mK$, $300\rm mK$. As can be seen from Figs. 11(a) and 11(b), in the absence of QOC, one can access the ground-state cooling condition $\langle\delta\hat{P}^{2}\rangle\approx\langle\delta\hat{Q}^{2}\rangle\approx 1/2$ until the temperature below $10\rm mK$. However, with increasing the temperature $T$, one has $\langle\delta\hat{P}^{2}\rangle\approx\langle\delta\hat{Q}^{2}\rangle>1/2$ so that $n_{\rm eff}>1$ [see Figs. 11(c) and 11(d)]. On the other hand, Figs. 11(e)-11(g) reveals that in the presence of QOC we have both $n_{\rm eff}<1$ and $\langle\delta\hat{P}^{2}\rangle\approx\langle\delta\hat{Q}^{2}\rangle\approx 1/2$ until the temperature $T=100\rm mK$. According to Fig. 11(h) we understand that as the temperature $T$ increases up to $300\rm mK$ the ground-state cooling requirement $n_{\rm eff}<1$ is still satisfied, though the energy equipartition breaks down.
@@
Figure 13: (Color online) Squeezing degree of the $Q$-quadrature of the MO vs the dimensionless microwave detuning $\Delta_{0w}/\Omega_{m}$ for (a) different values of the QOC and (b) different temperatures. In panel (a) the green-dotted, blue-dashed, and red-solid lines, respectively, correspond to $g_{2}=0$, $g_{2}/g_{1}=10^{-3}$, and $g_{2}/g_{1}=8\times 10^{-3}$, and the temperature is $T=1$mK. In panel (b) the blue-dashed, green-dotted, and red-solid lines, respectively, correspond to $T=1$mK, $T=50$mK, and $T=100$mK, and we have set $g_{2}/g_{1}=8\times 10^{-3}$. The shadowed pink region corresponds to squeezing beyond the 3 dB limit. The other parameters are the same as in Fig. 2.
@@
Based on numerical calculations, we find that the $P$-quadrature fluctuations cannot be less than $1/2$, while the $Q$-quadrature fluctuations can be less than $1/2$. Therefore, in what follows we will focus our discussion only on the squeezing of the position quadrature. In Fig. 12 we have plotted the squeezing degree of the $Q$-quadrature, $S_{Q}$, in dB as a function of $\Delta_{0w}/\Omega_{m}$ and $g_{2}/g_{1}$. As is seen, for the positive sign of QOC and in the vicinity of the microwave resonance frequency, $\Delta_{0w}/\Omega_{m}=0$, the 3-dB limit of squeezing can be beaten for the $Q$-quadrature of the MO so that it may be squeezed to high degrees (up to 7 dB). Nevertheless, mechanical squeezing is not possible for zero or negative values of QOC. In Fig. 12 we have shown the behavior of $S_{Q}$ versus the positive values of the ratio $g_{2}/g_{1}$ for $\Delta_{0w}/\Omega_{m}=0$. The figure shows that the mechanical squeezing ($S_{Q}>0$ dB) starts to appear at $g_{2}/g_{1}=0.001$, and strong mechanical squeezing($S_{Q}>$ 3dB) is achievable for $0.0230.0015<g_{2}/g_{1}<$ 0.023. We present the squeezing degree of the $Q$-quadrature with respect to $\Delta_{0w}/\Omega_{m}$ for three different non-negative values of $g_{2}/g_{1}$ in Fig. 13. As can be seen, in the absence of the QOC no squeezing is observed in the mechanical mode. However, for larger positive values of the QOC the beating of the standard squeezing limit of 3 dB is possible. Finally, we display in Fig. 13 the impact of the environment temperature $T$ on the squeezing degree of the $Q$ quadrature of the MO in the presence of QOC. Evidently, the degree of squeezing strongly decreases with increasing temperature.
@@ VI Conclusions
In summary, we have investigated the effects of quadratic coupling between mechanical and optical modes on the steady-state bipartite entanglements, mechanical ground-state cooling, and mechanical quadrature squeezing in a hybrid electro-optomechanical system. To do this, we assumed that the MO is linearly coupled to the microwave mode, while it is simultaneously coupled to the optical mode through both LOC and QOC. Our findings have revealed that the presence of even a weak QOC may lead to enhanced and thermally resistive bipartite entanglement between the subsystems. In particular, the stationary entanglement between the optical and microwave modes can be amplified by about 2.25 times around the temperature of 1mK, and it can be protected from the influence of thermal noise up to the ambient temperature of 90mK. By analyzing the ground-state cooling of the MO, we have found that in the optical and microwave red-detuned regime the mechanical ground-state cooling can be enhanced by about 1 order of magnitude by tuning the QOC parameter. Finally, we have concluded that for the positive sign of QOC and in the vicinity of the microwave resonance frequency, the 3-dB limit of squeezing can be surpassed up to 7dB for the $Q$-quadrature of the mechanical mode.
Appendix A Derivation of Hamiltonian
@@
Appendix B Stability conditions

### v2

@@ II.3 Stationary quantum fluctuations
[…]
Figure 11: (Color online)(a) Density plot of $S_{Q}$ vs $g_{2}/g_{1}$ and the normalized bare dimensionless detuning of the microwave cavity $\Delta_{0w}/\Omega_{m}$, and (b) plot of $S_{Q}$ vs $g_{2}/g_{1}$ for $\Delta_{0w}=0$. The shadowed pink region corresponds to squeezing beyond the 3 dB limit. Here we have set $\mathcal{P}=1mW$ and $\mathcal{P}_{w}=1mW$. The other parameters are the same as in Fig. 2.
It should be noted that the condition $n_{\rm eff}<1$ is not the only requirement for the ground-state cooling of the mechanical mode. In order to achieve the quantum ground-state cooling it is also necessary that both momentum and displacement variances tend to $\langle\delta\hat{P}^{2}\rangle\approx\langle\delta\hat{Q}^{2}\rangle\approx 1/2$ that is the energy equipartition should be satisfied in the optimal regime close to the ground state. Otherwise, the steady state of the system is not a thermal equilibrium state, and thus one cannot define an unambiguously effective temperature for the MO in this case genes20081 . We have examined this second requirement in Appendix D.
@@
Figure 12: (Color online) Squeezing degree of the $Q$-quadrature of the MO vs the dimensionless microwave detuning $\Delta_{0w}/\Omega_{m}$ for (a) different values of the QOC and (b) different temperatures. In panel (a) the green-dotted, blue-dashed, and red-solid lines, respectively, correspond to $g_{2}=0$, $g_{2}/g_{1}=10^{-3}$, and $g_{2}/g_{1}=4\times 10^{-3}$, and the temperature is $T=1$mK. In panel (b) the blue-dashed, green-dotted, and red-solid lines, respectively, correspond to $T=0.001$K, $T=0.05$K, and $T=0.115$K, and we have set $g_{2}/g_{1}=8\times 10^{-3}$. The shadowed pink region corresponds to squeezing beyond the 3 dB limit. Here we have set $\mathcal{P}=1mW$ and $\mathcal{P}_{w}=1mW$. The other parameters are the same as in Fig. 2.
@@
Based on numerical calculations, we find that the $P$-quadrature fluctuations cannot be less than $1/2$, while the $Q$-quadrature fluctuations can be less than $1/2$. Therefore, in what follows we will focus our discussion only on the squeezing of the position quadrature. In Fig. 11 we have plotted the squeezing degree of the $Q$-quadrature, $S_{Q}$, in dB as a function of $\Delta_{0w}/\Omega_{m}$ and $g_{2}/g_{1}$. As is seen, for the positive sign of QOC and in the vicinity of the microwave resonance frequency, $\Delta_{0w}/\Omega_{m}=0$, the 3-dB limit of squeezing can be beaten for the $Q$-quadrature of the MO so that it may be squeezed to high degrees (up to 15 dB). Nevertheless, mechanical squeezing is not possible for zero or negative values of QOC. In Fig. 11 we have shown the behavior of $S_{Q}$ versus the positive values of the ratio $g_{2}/g_{1}$ for $\Delta_{0w}/\Omega_{m}=0$. The figure shows that the mechanical squeezing ($S_{Q}>0$ dB) starts to appear at $g_{2}/g_{1}=0.001$, and strong mechanical squeezing($S_{Q}>$ 3dB) is achievable for $0.040.0015<g_{2}/g_{1}<$ 0.04. We present the squeezing degree of the $Q$-quadrature with respect to $\Delta_{0w}/\Omega_{m}$ for three different non-negative values of $g_{2}/g_{1}$ in Fig. 12. As can be seen, in the absence of the QOC no squeezing is observed in the mechanical mode. However, for larger positive values of the QOC the beating of the standard squeezing limit of 3 dB is possible. Finally, we display in Fig. 12 the impact of the environment temperature $T$ on the squeezing degree of the $Q$ quadrature of the MO in the presence of QOC. Evidently, the degree of squeezing strongly decreases with increasing temperature. However, as is seen, the quadrature squeezing may survive up to $T=0.115$K for $g_{2}=8\times 10^{-3}g_{1}$.
@@ VI Conclusions
To conclude, we have investigated the effects of quadratic coupling between mechanical and optical modes on the steady-state bipartite entanglements, mechanical ground-state cooling, and mechanical quadrature squeezing in a hybrid electro-optomechanical system. To do this, we assumed that the MO is linearly coupled to the microwave mode, while it is simultaneously coupled to the optical mode through both LOC and QOC. Here, we summarize the main advantageous features of the model under consideration: (i) In the presence of QOC the steady-state entanglement between optical and mechanical modes can be amplified by about 2 order of magnitude around the temperature of 1mK. In addition, the QOC can make the enhanced optical-mechanical entanglement robust against the thermal noise, such that it survives up to 0.1K. (ii) In the optical and microwave red-detuned regime, the mechanical ground-state cooling can be enhanced by about 1 order of magnitude through tuning the QOC parameter. (iii) In the presence of QOC with a positive sign, the 3-dB limit of mechanical squeezing can be surpassed up to 15dB in the vicinity of the microwave resonance frequency. The generated mechanical squeezing may survive up to the temperature $T=0.115$K .
Appendix A Derivation of Hamiltonian (2)
@@
Appendix B Linearized quantum Langevian equations
The linearized quantum Langevin equations for the fluctuation operators read as
$\delta\dot{\hat{Q}}=\Omega_{m}\delta\hat{P},$ (28b)
$\qquad-\gamma_{m}\delta P+\sqrt{2\gamma_{m}}\delta P_{in},$ (28c)
where $\tilde{G}=g_{1}+2g_{2}\langle\,Q\rangle_{ss}$ is the effective single-photon optomechanical coupling strength, modified by the QOC. Here, without loss of generality, we can take $\langle a\rangle_{ss}$ and $\langle a_{w}\rangle_{ss}$ to be real by appropriate choice of the phases of the driving fields $E_{d}$ and $E_{dw}$. By introducing the dimensionless optical-field quadratures fluctuations as $\delta\hat{Q}_{o}=\dfrac{\delta\hat{a}+\delta\hat{a}^{\dagger}}{2}$ and $\delta\hat{P}_{o}=\dfrac{\delta\hat{a}-\delta\hat{a}^{\dagger}}{2i}$, and also the dimensionless microwave-field quadratures fluctuations as $\delta\hat{Q}_{w}=\dfrac{\delta\hat{a}_{w}+\delta\hat{a}_{w}^{\dagger}}{2}$ and $\delta\hat{P}_{w}=\dfrac{\delta\hat{a}_{w}-\delta\hat{a}_{w}^{\dagger}}{2i}$, the set of Eqs. (28) can be expressed in the compact matrix form
where
$\delta\hat{u}=(\delta\hat{Q}_{o},\delta\hat{P}_{o},\delta\hat{Q},\delta\hat{P},\delta\hat{Q}_{w},\delta\hat{P}_{w})^{T}$ (30)
is the vector of continuous-variable fluctuations and A is the $6\times 6$ drift matrix which is given by
and
$\delta\hat{n}(t)=$ $(\sqrt{2\kappa}\delta\hat{Q}_{in,o},\sqrt{2\kappa}\delta\hat{P}_{in,o},0,\sqrt{2\gamma_{m}}\delta\hat{P}_{in},\sqrt{2\kappa_{w}}\delta\hat{Q}_{in,w}$ (32)
$,\sqrt{2\kappa_{w}}\delta\hat{P}_{in,w})^{T}$
represents the corresponding vector of noises.
Appendix C Stability conditions
@@
Appendix D Ground-state cooling and energy equipartition
Figure 13: (Color online) Plots of effective mean phonon number $n_{\rm eff}$ (on logarithmic scale) and the variances $\langle\delta\hat{Q}^{2}\rangle-3/4$ and $\langle\hat{P}^{2}\rangle-3/4$ versus the normalized detuning $\Delta_{0w}/\Omega_{m}$ for $g_{2}=0$ (panels a and b) and $g_{2}=8\times 10^{-3}g_{1}$ (panels c and d), and for different values of the environment temperature $T$. Other parameters are the same as in Fig. 2.
We have examined the condition of energy equipartition $\langle\delta\hat{P}^{2}\rangle\approx\langle\delta\hat{Q}^{2}\rangle\approx 1/2$ in Fig. 13, where we have plotted the effective mean phonon number $n_{\rm eff}$ (on logarithmic scale) and the variances $\langle\delta\hat{Q}^{2}\rangle-3/4$ and $\langle\delta\hat{P}^{2}\rangle-3/4$ versus $\Delta_{0w}/\Omega_{m}$ for $g_{2}=0$ (panels a and b) and $g_{2}=8\times 10^{-3}g_{1}$ (panels c and d), and for different values of temperature $T=0.1\rm K$ and $0.3\rm K$. As can be seen, in the absence of QOC, energy equipartition breaks down, i.e., one has $\langle\delta\hat{P}^{2}\rangle\approx\langle\delta\hat{Q}^{2}\rangle>1/2$ so that $n_{\rm eff}>1$ [see Figs. 13(a) and 13(b)]. On the other hand, in the presence of QOC we have both $n_{\rm eff}<1$ and $\langle\delta\hat{P}^{2}\rangle\approx\langle\delta\hat{Q}^{2}\rangle\approx 1/2$ until the temperature $T=0.1\rm K$. According to Fig. 13(d) we understand that as the temperature $T$ increases up to $0.3\rm K$ the ground-state cooling requirement $n_{\rm eff}<1$ is still satisfied, though the energy equipartition breaks down.